### v1

# DeepQuestion: Systematic Generation of Real-World Challenges for Evaluating LLMs Performance

Abstract
LLMs often excel on standard benchmarks but falter on real-world tasks. We introduce DeepQuestion, a scalable automated framework that augments existing datasets based on Bloom’s taxonomy and creates novel questions that trace original solution paths to probe evaluative and creative skills. Extensive experiments across ten open-source and proprietary models—covering both general-purpose and reasoning LLMs—reveal substantial performance drops (even up to 70% accuracy loss) on higher-order tasks, underscoring persistent gaps in deep reasoning. Our work highlights the need for cognitively diverse benchmarks to advance LLM progress. DeepQuestion and related datasets will be released upon acceptance of the paper.
DeepQuestion: Systematic Generation of Real-World Challenges for Evaluating LLMs Performance
Ali Khoramfar, Ali Ramezani, Mohammad Mahdi Mohajeri Mohammad Javad Dousti, Majid Nili Ahmadabadi, Heshaam Faili Department of Electrical and Computer Engineering University of Tehran {khoramfar, ali.ramezani.96, mehdimohajeri, mjdousti, mnili, hfaili}@ut.ac.ir
1 Introduction
Large Language Models (LLMs) have shown strong performance on a variety of standardized benchmarks, such as GSM8K, MATH, MMLU, MMLU-Pro, and GPQA Achiam et al. (2023); Team et al. (2023); Grattafiori et al. (2024); Team et al. (2025); Cobbe et al. (2021); Hendrycks et al. (2021b, a); Wang et al. (2024); Rein et al. (2024). However, these benchmarks are highly structured and may not capture the complexity and variability found in real-world problems Arias-Duart et al. (2025); Myrzakhan et al. (2024).
Recent work GSM-Symbolic Mirzadeh et al. (2025) shows that adding extra numerical values and irrelevant information to benchmark questions reduces LLM performance, highlighting their difficulty with extraneous content. Since real-world tasks often include such irrelevant details, this limitation raises concerns about the practical robustness of LLMs.
Proof or Bluff Petrov et al. (2025) finds that while LLMs perform well on older math contests, they struggle with the 2025 USA Mathematical Olympiad, with top models like Gemini-2.5-Pro scoring below 25. This indicates a reliance on training data patterns rather than genuine reasoning abilities.
Alaa et al. (2025) report that while LLMs excel on medical exams like the USMLE, these tests lack construct validity and do not reflect real clinical complexity. Thus, strong benchmark results do not guarantee clinical effectiveness. Scenario-based questions that assess reasoning, rather than recall, are needed for a more accurate evaluation of higher-order cognition Chauhan et al. (2025).
Learning science offers frameworks for understanding learning levels, Bloom’s Taxonomy Forehand (2010) is a widely recognized framework outlining six cognitive skill levels, from basic recall to higher-order reasoning Smith (2008); Leung (2000). LLMs often excel at recall tasks but struggle with novel problems Petrov et al. (2025), indicating limited generalization beyond memorized data.
Building on Bloom’s Taxonomy, we introduce the DeepQuestion framework to create new benchmarks from existing datasets. It generates (1) real-world scenarios to test application and analysis, and (2) instructions for crafting similar questions to assess creation, evaluation, and instruction-following, enabling deeper evaluation of LLMs’ knowledge. Our main contributions are:
  • Proposed DeepQuestion, a framework for generating questions from existing datasets for better LLM evaluation based on Bloom’s taxonomy.
  • Introduced the DeepQuestion dataset, created using our framework.
  • Conducted a comprehensive evaluation of LLMs with the DeepQuestion dataset, highlighting their knowledge limitations across Bloom’s taxonomy.
The remainder of this paper is organized as follows: Section 2 introduces the DeepQuestion framework, Section 3 details our experimental setup, and the final sections present our results and conclusions.
2 Method
This study is guided by a key research question: Can existing benchmarks be deepened to reflect real-world tasks more accurately?
To address this question, we draw on educational learning science, especially Bloom’s taxonomy, which organizes cognitive skills into six levels. Prior works show that most benchmark questions focus on the first two levels—remembering and understanding Alaa et al. (2025); Chauhan et al. (2025). We hypothesize that deepening these questions along Bloom’s hierarchy could better engage higher-order reasoning and reflect real-world complexity. Therefore, we introduce the DeepQuestion framework, which generates more insightful evaluation questions through two methods: (1) question-to-scenario (Q2S) and (2) question-to-instruction (Q2I) generation. The details of these methods, as well as the prompt generation pipeline that automates their application to existing benchmarks, are described in the rest of this section.
2.1 Question-to-Scenario Generation (Q2S)
To deepen questions, Q2S targets the third level of Bloom’s taxonomy, namely apply, which emphasizes on using knowledge in practical contexts. Each original question is transformed into a scenario-based version embedding the core problem within a realistic narrative, often including extraneous details and some distractions to simulate real-world complexity. For example, a basic physics question asking for the position-time equation of an object, is reframed within a scenario involving an environmental drone, where irrelevant elements like weather acted as distractors. The core question data and required reasoning remain unchanged. Q2S employs an automated prompt-based pipeline powered by an LLM (see Subsection 2.3 for more details on the pipeline). The output is a benchmark of scenario-driven questions that better reflect authentic problem-solving contexts.
2.2 Question-to-Instruction Generation (Q2I)
Encouraged by the initial result analysis at the Bloom’s apply level with Q2S, we extend our approach to the higher levels of evaluate and creation. At these levels, learners are expected to assess information critically and produce novel outputs. We hypothesize that individuals or models with deep conceptual understanding are not only able to solve problems but also to design meaningful questions. This informs our second experimental setup, which reverses the direction of the task: instead of solving a question, the model is instructed to design one.
For each original dataset question, Q2I constructs an instruction designed to prompt a language model to generate a new question that preserves the same topic and solution path as the original. This instruction is formulated in a way that ensures a different LLM, when given the instruction, would also produce a question aligned with the original in both topic and reasoning process. For example, if the original question in Subsection 2.1 (position-time equation of an object) yields the solution $x=3t-4$, the instruction would ask the model to design a physics problem involving uniform linear motion whose solution would be exactly that equation. Designing questions based on instruction implicitly tests three levels of the model’s reasoning: conceptual understanding of the domain and relevant equations, the ability to define appropriate variables and their interrelations, and the selection of values that yield the target solution.
Q2I employs an automated prompt-based pipeline similar to Q2S; however, the output of Q2I is the set of instructions for question generation. The next subsection details the pipeline.
Figure 1: Overview of DeepQuestion framework. It begins with the selection of a random batch of questions and answers. Then, by conversation between the prompt generator and the prompt evaluator LLMs, the deep-question prompt is generated. The question generator LLM with the deep-question prompt converts each question and answer pair to the deep question and answer.
2.3 Prompt Generation Pipeline
Since transforming each benchmark into its Q2S and Q2I variants requires carefully crafted prompts tailored to the specific domain and style of that benchmark, a key component of the DeepQuestion framework is a prompt generation pipeline which automates the creation of these task-specific prompts and is shown in Figure 1. This pipeline eliminates the dependency on human expertise for prompt design and enables scalable adaptation to different benchmarks.
The pipeline operates through an iterative dialogue between two LLMs: a prompt generator and a prompt evaluator. Initially, a randomly selected batch of questions and their answers are sampled from the source benchmark, along with a high-level goal description, are provided to the prompt generator LLM. The goal is to produce a prompt that can transform similar input questions into scenario-based questions in Q2S or instruction prompts in Q2I.
The generated prompt is then passed to the evaluator LLM, which assesses it based on predefined evaluation criteria. The evaluator assigns a numerical score ranging from 0 to 10 and provides qualitative feedback describing strengths and weaknesses of the prompt.
If the score surpasses a threshold (e.g., 8), the prompt is accepted for use in the corresponding generation task. Otherwise, the prompt generator revises the prompt using the evaluator’s feedback, and the process repeats iteratively until a satisfactory prompt is produced.
This automatic prompt optimization is described in Algorithm 1 in the Appendix Section. By leveraging this pipeline, the DeepQuestion framework achieves robust prompt design tailored to the specific domain and style of each benchmark, enabling effective question transformation without manual intervention.
3 Experiment Setup
We constructed the DeepQuestion benchmark by applying our framework to 60 randomly selected GSM8K questions (in English) and 60 physics questions from the Iranian University Entrance Exam (in Persian). While our method aims to minimize reliance on expert intervention, these datasets were chosen due to the familiarity of the authors to math and physics subjects. The generated deep questions were manually reviewed for quality and accuracy, while we evaluated the framework by comparing them to the original questions.
(a) Original vs. Q2I in GSM8K
(b) Original vs. Q2I in Physics
Figure 2: Evaluation of different LLMs in original and instruction-based question
For evaluation, we utilized, Gemini-2, GPT-4.1, Llama-3.1 Grattafiori et al. (2024), Deepseek-V3 Liu et al. (2024), Deepseek-R1 Guo et al. (2025), O4-mini, Gemma3 Team et al. (2025), Phi4 Abdin et al. (2024), and Qwen3 Yang et al. (2025). To ensure reproducibility, all models were run with a temperature setting of zero. Furthermore, within the framework itself, we employed Gemini 2.5 Pro in prompt generator LLM, prompt evaluator LLM, and question generator LLM.
In the Q2S setting, the model generates a solution to a given scenario-question, and correctness is easily verified due to well-defined answers in math and physics problems in our setup.
In the Q2I setting, the model generates questions from instructions. Evaluation consisted of two steps: (1) Answerability Check: verifying whether a strong model (O4-mini) and the model itself can correctly solve the generated questions, indicating questions validity. (2) LLM-as-Judge: since answerability alone does not capture question quality, a powerful language model (O4-mini) was employed for direct qualitative evaluation based on expert-defined criteria.
4 Results
4.1 General-Purpose LLMs
First, we evaluated general-purpose LLMs on question generation using Q2S. Moderate performance loss in Q2S (see Appendix C) motivated us to design the Q2I experiment, where models showed a substantial accuracy drop for tasks demanding deeper understanding. No model surpassed 38% accuracy on instruction-based question generation for GSM8K or physics (see Figure 2), despite over 95% on original questions. These results support the use of Bloom’s taxonomy as an evaluation framework and reveal the limitations of general-purpose LLMs on higher-order cognitive tasks.
4.2 Reasoning LLMs
Reasoning models maintain relatively stable performance in the Q2S setting, indicating proficiency at the Bloom’s apply level (Appendix C). However, in the Q2I setting, accuracy drops by 8–30% (Figure 2), exposing limitations at higher cognitive levels. For instance, Deepseek-R1’s performance decreases by 30% on GSM8K and 10% on physics, while Qwen3-32b drops by 11% on both. Although reasoning models outperform general-purpose models, substantial gaps remain in achieving deep, comprehensive understanding.
Figure 3: Win rate of original against Q2I physics questions for O4-mini.
4.3 Question Quality Analysis
While reasoning models can generate questions that follow instructions and maintain topical consistency, it is crucial to assess their quality beyond surface-level compliance. To this end, we use a preference-based evaluation with O4-mini as a judge, directly comparing generated and original questions across five criteria: reasoning demand, numerical quality, physical realism, clarity and brevity, and solution spoiling (see Appendix B for explanation of each criterion). This approach addresses concerns about the model’s ability to assess quality by focusing on direct comparison.
As shown in Figure 3 and Appendix D, generated physics questions still lag behind the originals, reflecting ongoing challenges for reasoning models. In contrast, generated GSM8K questions are more comparable to the originals, likely due to the dataset’s simplicity. These results confirm the effectiveness of our framework in increasing question depth without introducing artificial difficulty. However, a significant quality gap persists, highlighting the ongoing challenge of generating high-quality questions, even for advanced reasoning models.
5 Conclusion
Despite strong performance on standard benchmarks, LLMs struggle with tasks requiring deeper and real-world reasoning. We introduced DeepQuestion, a scalable automated framework which enhances existing datasets based on Bloom’s taxonomy. Our method exposes reasoning gaps often missed by conventional evaluations. Applied to GSM8K and physics datasets, our framework reveals sharp accuracy drops, up to 70%, on complex reasoning tasks across a wide range of models. These findings highlight serious limitations in current LLMs and underscore the need for cognitively rich benchmarks.
Limitations
Although our proposed framework is capable of handling open-ended questions, in this study we restrict our focus to questions with definite correct answers. To enable expert human evaluation, we concentrate on two benchmarks related to mathematics and physics. Nonetheless, our framework is not specifically tailored to these benchmarks and can be extended to other domains as well.
Appendix A Algorithm
Algorithm 1 describes the iterative dual-LLM pipeline for generating and evaluating prompts that transform an existing QA dataset into deep, scenario- and instruction-based questions.
Input: QA dataset $B=\{(q_{i},a_{i})\}$
Output: Deeper Question Benchmark: DeepQuestion
1
[…]
4goal_description $\leftarrow$ Description of the intended purpose of generated prompts;
5evaluation_criteria $\leftarrow$ Instructions for how the LLM should assess generated prompt quality;
6Sample a random subset $B_{s}\subset B$;
7while $score<8$ do
8       $prompt\leftarrow\texttt{LLM}(goal\_description,B_{s},feedback)$;
9      $(score,feedback)\leftarrow\texttt{LLM}(evaluation\_criteria,prompt)$;
10
11DeepQuestion $\leftarrow$ empty list;
12foreach $(q,a)\in B$ do
13       $q^{\prime}\leftarrow\texttt{LLM}(prompt,q,a)$; DeepQuestions.append$(q^{\prime},a)$;
Algorithm 1 DeepQuestion Framework Algorithm
Appendix B Prompts
Below are the core prompt templates used in the DEEPQUESTION framework, including the Q2S scenario-generation and Q2I instruction-generation prompts, as well as the system-level prompts for prompt evaluation and question generation.
System Prompt for Questions Quality Evaluation in Q2I You are an expert evaluator of high school physics problems. You will compare two problems, A and B, which lead to the same numeric answer (provided only for verification). Your job is to judge which problem is better, based on meaningful teaching and assessment value — not on how easy or flashy it appears. Before making your judgment, consider: 1. What physical reasoning each problem requires. 2. How the numbers and context affect realism and difficulty. 3. How clearly and economically the problem is written. 4. Whether the problem gives away the solution or answer inappropriately. — ### PROBLEMS A: {original_q} B: {generated_q} target_answer = {target_answer} — ### CRITERIA & GUIDANCE 1. Reasoning Demand - Definition: How much actual thinking and decision-making the problem requires. - Better if: The solver must interpret the situation, choose a method, or justify an approach — not just apply a memorized formula. - Worse if: The solution is a single plug-in step with no real choices or insight required. 2. Numerical Quality - Definition: Thoughtfulness and realism of the numbers used. - Better if: Values are plausible and require genuine calculation; they are not trivial. - Worse if: The numbers oversimplify the math or seem artificially chosen just to make the answer neat (e.g., all variables are $1$ or 00). 3. Physical Realism - Definition: Whether the scenario is physically plausible and internally consistent. - Better if: Units, magnitudes, and assumptions make sense; any simplifications (e.g., ignoring friction) are clearly stated. - Worse if: The context contradicts basic physics or feels contrived or incoherent. 4. Clarity & Brevity - Definition: How concise and understandable the problem is. - Better if: The question is short, clear, and includes only relevant information. - Worse if: It contains unnecessary explanation, formula reminders, solution hints, or redundant language. 5. Solution Spoiling - Definition: Whether the problem gives away the answer or guides through steps. - Better if: The question asks for a result without hinting at how to obtain it. - Worse if: The final answer is stated, or if key solution steps are included (e.g., "first calculate time, then…"). — ### DECISION RULES - For each criterion, assign the better problem: A, B, or E (Equal). - Use E only if the difference is truly negligible — avoid it unless absolutely necessary. - Be decisive and clear. — ### OUTPUT Produce exactly this JSON (no extra text): REMEMBER: use E only when A and B are nearly indistinguishable. {{ "Reasoning_Demand": "<A|B|E>", "Numerical_Quality": "<A|B|E>", "Physical_Realism": "<A|B|E>", "Clarity & Brevity": "<A|B|E>", "Solution_Spoiling": "<A|B|E>" }}
Deep-Question Prompt Q2S for Physics Questions My goal is to design questions based on Bloom’s *Apply* level. I will give you a question, and your task is to create a scenario in the form of a narrative that includes redundant material not related to solving the question. Ultimately, the explicit question I provide must be solved to find the answer, but the student must infer this question from the narrative. The narrative should include a real-world story and irrelevant numbers that are ultimately simplified to the original problem. * You should never give any hints about which information is necessary and which is irrelevant. * The story should not reference the formulas or concepts required to solve the question. It should put the solver in a situation where they must apply their knowledge independently. * All formulas and numbers should be presented in LaTeX format. * All your answers must be in Persian.
Deep-Question Prompt in Q2I for Physics Questions My goal is to design questions based on Bloom’s *Assessment* level. I will provide you with a question. Your task is to write a simple, non-detailed instruction that students will use to design a new question**—one that leads to **exactly the same final answer as the original question. The aim is to assess whether students have a deep enough understanding of the concept to create a question that results in a specific outcome. ### The following conditions must be met: * The instruction should not explicitly suggest the use of any formula. * The instruction should not include specific details about question design (e.g., what parameters or values are needed). * The final output (i.e., the answer) must be clearly stated in the instruction and must match the original question’s final answer. * Since the instruction is given independently of the original question, it must include all the context/data required. * The instruction should not provide direct guidance and must resemble an exam-style prompt. * The designed question must not be multi-part; it should have one clear requirement that is also explicitly stated. * All formulas and numbers must be presented in LaTeX format. * The instruction must be written in Persian. ### Example: Design a question that, using the concepts of motion, results in a space-time equation with the final output $x=3t-4$
Prompt for Evaluation of Answers You are a helpful assistant. You will receive a physics question. Your task is to answer it in Persian using English numerals and LaTeX format for all formulas and numerical expressions. Solve the problem step by step (chain-of-thought reasoning) before providing the final answer. Ensure clarity and logical reasoning throughout your explanation. If the final answer is a number or formula, repeat it exactly at the end, after the following line: ### answer: Only the final answer should appear after the "### answer:" line, with no extra text, punctuation, or explanation. Do not include anything after the "### answer:" section.
System Prompt Question Generator in Q2I "You are a helpful assistant. Design a physics question based on the input instruction and consider the following: * All formulas and numbers should be in English and LaTeX format. * All your answers must be in Persian. * Just design the question; do not answer it. * Do not include the final answer in your question. * The designed question should not be multi-part; it should ultimately have one requirement, and that requirement must also be specified in the instruction given to you. * Remember you are NOT ALLOWED to give the final answer explicitly in the question. * You must think step by step before generating the question. Put your reasoning process between <think> and </think> tags. * The final designed question in Persian must be written immediately after the </think> tag."
System Prompt for Prompt Generator in Physics Questions You are a prompt engineer tasked with writing a Scenario Generation Instruction (SGI). This SGI will be used to guide a language model in generating original, scenario-based physics questions. Your SGI must instruct the model to: Create questions grounded in realistic or imaginative physical scenarios. Ensure each question tests the same underlying physics principles as a provided set of example questions. Avoid copying or making superficial edits to the example questions. Your final output should be a well-structured, standalone prompt that clearly communicates these instructions to the model. Output only the SGI.
System Prompt for Prompt Evaluator in Physics Questions You are a helpful assistant assigned to evaluate a Scenario Generation Instruction (SGI), which is designed to guide another model in generating high-quality, scenario-based physics questions based on a sample set of physics Q&A. Your task is to: Evaluate the SGI according to these criteria: Clarity: Are the instructions understandable and unambiguous? Relevance: Does it stay focused on the concepts represented in the example questions? Instructional Effectiveness: Will it likely result in creative and accurate scenario-based questions? Creative Direction: Does it provide sufficient guidance to inspire realistic or imaginative scenarios rooted in physics? Provide Feedback: Suggest specific improvements to strengthen the SGI. Score the SGI: Assign a numerical score between 00 (very poor) and $10$ (excellent) based on overall quality. Your output must include: A brief evaluation summary (strengths and weaknesses). Clear, actionable suggestions for improvement. A final numeric score between 00 and $10$.
Appendix C Q2S Results
These charts show each model’s accuracy on the original benchmark questions compared to their scenario-based (Q2S) versions, illustrating the performance drop when questions are embedded in realistic narratives.
Figure 4: Result of Original vs.Q2S in Physics
Figure 5: Result of Original vs. Q2S in GSM8K
Appendix D Quality of Generated Question
These plots display, for each of five quality criteria, the percentage of cases where original questions “win” over generated ones, highlighting which aspects (e.g., reasoning demand, clarity) most favor the originals.
Figure 6: Win rate original against Q2I for O4-mini in GSM8K questions
Figure 7: Win rate original against Q2I for Deepseek-R1 in physics questions
Figure 8: Win rate original against Q2I for Deepseek-R1 in GSM8K questions
Figure 9: Win rate original against Q2I for Qwen3-32b in physics questions
Figure 10: Win rate original against Q2I for Qwen3-32b in GSM8K questions